OpenStudy (anonymous):

the formula for converting degrees Fahrenheit (F) to degrees Celsius (C) is C = 5/9 (F - 32). find C for F = 41

OpenStudy (***[isuru]***):

C = 5/9 ( F - 32) = 5/9( 41 - 32) = 5 got it :) ?

OpenStudy (anonymous):

got it :)